OpenStudy (anonymous):

Z a total of 65 coins, all of which are either dimes or nickels. The total value of the coins is $5.95. Find the number of each type of coin. @Mathematics

OpenStudy (anonymous):

Dime=65-x Nickel=x .10(65-x)+.05(x)=5.95 6.5-.10x+.05x=5.95 -.05x=.55 x=11 There are 11 Nickels And plug the x value into the "Nickel equation" 65-11=54 There are 54 Dimes